OpenStudy (anonymous):

if cosx 5/13 and sinx < 0, find sin(x/2)

OpenStudy (anonymous):

sorry, its cosx = 5/13

OpenStudy (anonymous):

Shall we work together on this?

OpenStudy (anonymous):

sure, why not? :D

OpenStudy (anonymous):

Awesome! Let's do this. Let me warn you: this will take a little bit of time to do. http://www.twiddla.com/584559 Let's go here because trig is better done on "paper".

OpenStudy (anonymous):

This is a 5 12 13 Pythag triangle and Sin is negative in 3 and 4 quadrant....

OpenStudy (anonymous):

my book's answer is actually negative... which is why im asking this question

OpenStudy (anonymous):

Wait, negative or positive?

OpenStudy (anonymous):

Should be negative, as it's in the 4th quadrant

OpenStudy (anonymous):

You could use the identity \[cos(x)=1-2sin^2\left(\frac{x}{2}\right)\]

OpenStudy (anonymous):

Or draw the appropriate triangle and just bisect the angle..

OpenStudy (anonymous):

this is what you do if you want easy way. look up "half angle formula" for sine, which says \[\sin(\frac{x}{2})=\pm\sqrt{\frac{1-\cos(x)}{2}}\]

OpenStudy (anonymous):

replace \[\cos(x)\] by \[\frac{5}{13}\] because that is what you are told. you are also told that you want the negative one. substitute and be done

OpenStudy (anonymous):

get \[-\sqrt{\frac{1-\frac{5}{13}}{2}}\] then some arithmetic to get \[-\sqrt{\frac{\frac{8}{13}}{2}}=-\sqrt{\frac{4}{13}}=-\frac{2}{\sqrt{13}}\]\]

OpenStudy (anonymous):

@Satellite: We got that too, but his/her book says it's the same thing in positive. So the book is wrong?

OpenStudy (anonymous):

oh now i have to think. hold on while i turn on my battery

OpenStudy (anonymous):

ok it helps if i read. it says \[\sin(x)<0\] not \[\sin(\frac{x}{2})<0\]

OpenStudy (anonymous):

cosine is positive, sine is negative. you are in quadrant IV (don't you love roman numerals?)

OpenStudy (anonymous):

so when you take half of x you will be somewhere up in quadrant II and sine will be positive. so it is positive. sorry

OpenStudy (anonymous):

answer is the positive root not negative one. my fault.

OpenStudy (anonymous):

i dont really understand the signs...

OpenStudy (anonymous):

it starts at quadrant IV, so shouldn't the formula be\[\sin (x/2) = -(\sqrt{1-cosx}/2)\]?